Test the degree of association or significant relationship between the teaching
performance and management performance of the faculty members of school using the
Pearson product moment method at 0.05 level of significance. Give the decision.
Teachers Teaching
Performance
Management
Performance
A 3.6 3.8
B 4.4 3.9
C 4.1 2.0
D 4.0 2.8
E 4.0 2.9
F 3.4 3.0
G 2.8 4.5
H 4.9 3.3
I 3.6 2.9
J 3.0 3.0
K 3.5 3.8
L 3.0 2.7
M 4.3 3.2
N 3.2 3.3
O 4.3 3.4
P 3.4 4.1
Q 3.8 4.3
R 3.9 4.4
S 4.6 3.8
T 3.5 3.3
1
Expert's answer
2020-08-03T19:15:40-0400

Solution


r=(n(xy)(x)(y))([nx2(x)2][ny2(y)2])r= \frac {(n(∑xy)-(∑x)(∑y))}{\sqrt{([n∑x^2-(∑x)^2][n∑y^2-(∑y)^2]})}


xy=257.06∑xy=257.06

x=75.3∑x= 75.3

y=68.4∑y= 68.4

x2=289.59∑x^2 = 289.59

y2=241.86∑y^2 = 241.86

n=20n=20


r=(20257.0675.368.4)(([20289.5975.32][20241.8668.42]))r= \frac {(20*257.06-75.3*68.4)}{(\sqrt {([20*289.59- 75.3^2 ][20*241.86- 68.4^2])} )}

r=0.06707r= -0.06707


The test statistic:


t=(r(n2))(1r2)t= \frac {(r \sqrt{(n-2)})}{\sqrt{(1- r^2 )}}

t=(0.06707(202)(0.5)(1(0.06707)2)t= \frac{(-0.06707 * (20-2)^(0.5)}{\sqrt{(1- (-0.06707)^2 })}

=0.2852= -0.2852


The p-value given t=0.2852t = -0.2852 and n2=202=18n-2=20-2=18 degrees of freedom:

pvalue=0.7789p value=0.7789


Answer: t value = - 0.0670

                The correlation is not significant since p value (0.7789) > significant level (0.05)